Write down the dimensions of the following quantities: work, impulse, density, coefficient of viscosity, power, modulus of elasticity

1
Expert's answer
2021-12-30T11:53:54-0500

(a) By the definition of work, we have:


"W=Fd,"

here, "F" is the force, "d" is the displacement.

Then, we can write the dimensions of work as follows:


"W=[\dfrac{ML}{T^2}]\times [L]=[ML^2T^{-2}]."

(b) By the definition of the impulse, we have:


"J=Ft,"

here, "F" is the force, "t" is the time.

Then, we can write the dimensions of impulse as follows:


"J=[\dfrac{ML}{T^2}]\times[T]=[MLT^{-1}]."

(c) By the definition of the density, we have:


"\rho=\dfrac{m}{V},"

here, "m" is the mass, "V" is the volume.

Then, we can write the dimensions of density as follows:


"\rho=[\dfrac{M}{L^3}]=[ML^{-3}]."

(d) By the definition of the coefficient of viscosity, we have:


"\eta=\dfrac{F}{A\dfrac{dv}{dx}},"


here, "F" is the force, "A" is the area, "\dfrac{dv}{dx}" is the velocity gradient.

Then, we can write the dimensions of coefficient of viscosity as follows:


"\eta=[\dfrac{ML}{T^{2}}]\times[L^{-2}]\times[(\dfrac{LT^{-1}}{L})^{-1}]=[ML^{-1}T^{-1}]."


(e) By the definition of the power, we have:


"P=\dfrac{W}{t}=\dfrac{Fd}{t},"

here, "W" is work, "F" is the force, "d" is the displacement, "t" is the time.

Then, we can write the dimensions of power as follows:


"P=[\dfrac{ML}{T^2}]\times[L]\times[T^{-1}]=[ML^2T^{-3}]."

(f) By the definition of modulus of elasticity, we have:


"Y=\dfrac{stress}{strain}=\dfrac{FL}{A\Delta L},"

here, "F" is the force, "L" is the original length of the object, "A" is the cross-sectional area, "\Delta L" is the change in length of the object.

Then, we can write the dimensions of modulus of elasticity as follows:


"Y=\dfrac{[MLT^{-2}]\times[L]}{[L^2]\times[L]}=[ML^{-1}T^{-2}]."